# maximum a posteriori

### v1

@@ 3.1 MAP Estimator via BOLT
The proof is given in the supplementary file. ∎
Remark 3 (BOLT Bias and Variance).
Let $U(x)$ denote the likelihood ratio $U(x)=P_{1}(x)/P_{2}(x)$ and let $\widehat{U}$ be its plug-in estimator (see the supplementary file for construction details). Assume $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$. Then the BOLT plug-in Bayes-error estimator satisfies
$\bigl|\mathbb{E}\big[\widehat{\varepsilon}_{\mathrm{BOLT}}\big]-\varepsilon_{\mathrm{Bayes}}\bigr|$ $\;\leq\;q_{1}\,\varepsilon_{0}+\mathcal{O}\!\big(M^{-1/2}\big),$ (12)
$\mathrm{Var}\!\big(\widehat{\varepsilon}_{\mathrm{BOLT}}\big)$ $\;=\;\mathcal{O}\!\big(M^{-1}\big),$ (13)
where $M$ is the number of samples used in the plug-in estimator. Additional details are given in the supplementary file.
Theorem 2 gives an intuitive interpretation to the BOLT loss. It indicates that using BOLT, the model learns to mimic the Bayesian optimal classifier. From the computational point of view, using BOLT leads to gradients that guide the model to approximate the MAP estimator. This finding enables us to use the BOLT loss to train the discriminator of a GAN model. We illustrate this approach in the sequel.
BOLT-GAN
Let $s_{\theta}:\mathcal{X}\to\mathbb{R}$ denote a model which computes a raw score for input $x$, and let $h_{\theta}$ be a negative sigmoid activation of its output, i.e.
$h_{\theta}(x)=-\sigma(s_{\theta}(x)).$ (14)
From Theorem 2, one can conclude that in this case, the model $h_{\theta}$ estimates the Bayesian optimal discriminator.
Following a similar approach as in vanilla GAN, we can consider sampling from data and generator distributions conditionally with with $P_{1}=P_{\rm data}$ and $P_{2}=P_{\rm g}$ using a prior $\pi=P(\lambda=1)$ for some $\pi\in(0,1)$, where $\lambda=1$ and $\lambda=2$ correspond to real and fake samples, respectively. Noting that the generator samples are given by $g_{\phi}(Z)$ for latent samples $Z\sim P_{Z}$, the expected BOLT loss of the discriminator is given by
$R_{\text{BOLT}}(\phi,\theta)=-\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (15)
with $\mathcal{L}_{\text{BOLT}}(\phi,\theta)$ being
$\mathcal{L}_{\text{BOLT}}(\phi,\theta)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h_{\theta}(X)]$
$(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (16)
Given this risk function, the generator and discriminator are jointly trained by maximizing the minimum loss. This is equivalent to solving the following min-max game:
$\min_{\phi}\max_{\theta}\mathcal{L}_{\text{BOLT}}(\phi,\theta).$ (17)
Considering Theorem 1, one can see that the inner optimization minimizes the Bayes error upper bound. [^4] Note that $\mathcal{L}_{\text{BOLT}}=-R_{\text{BOLT}}$.
Stability of BOLT-GAN
The learning objective in BOLT-GAN is similar to that of WGAN, with two main differences: ($i$) BOLT-GAN considers an arbitrary sampling prior $\pi$, as in vanilla GAN, and ($ii$) it does not impose any Lipschitz continuity restriction on the discriminator model. The discussions in (Arjovsky et al.,, 2017; Gulrajani et al.,, 2017; Miyato et al.,, 2018) suggest that the latter difference can lead to lack of stability, when we use BOLT-GAN in this vanilla form. Our numerical experiments confirm the validity of this conjecture. To understand the reason behind this behavior, we study the convergence of the vanilla BOLT-GAN in the sequel. Our result indicates that BOLT-GAN implicitly minimizes the total variation (TV), which is a strict notion of distance. This finding enables us to modify the BOLT-GAN learning objective to a stable form.
3.2 Convergence of BOLT-GAN
Consider the general learning objective
$\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$ (18)
$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$
which is defined for any measurable $h:\mathcal{X}\mapsto[0,1]$ and measurable $g:\mathcal{Z}\mapsto\mathcal{X}$ that maps that latent space $\mathcal{Z}$ to the data space. $\mathcal{L}_{\text{BG}}^{(\pi)}$ reduces to $\mathcal{L}_{\text{BOLT}}$, when we replace $g$ and $h$ are parameterized models $g_{\theta}$ and $h_{\theta}$.
@@ 3.2 Convergence of BOLT-GAN
Our goal is to understand the metric, under which the generator distribution $P_{\rm g}$ converges to $P_{\rm data}$, when the min-max game with objective $\mathcal{L}_{\text{BG}}^{(\pi)}$ is solved. To this end, let us first define total variation (TV).
@@
The total variation between two distributions $P$ and $Q$ defined on a measurable $\mathcal{X}$ is defined as
$TV(P,Q)=\sup_{A\subseteq\mathcal{X}}|P(A)-Q(A)|,$ (19)
with the supremum taken over all events $A$ in the $\sigma$-field over $\mathcal{X}$.
The convergence of TV to zero implies for convergence of other distances, such as including JS and EM. The following lemma gives an alternative representation for TV.
Lemma 1 (Dudley, 2002a ).
@@
$TV(P,Q)=\sup_{f:\mathcal{X}\to[0,1]}\int f\,\mathrm{d}(P-Q),$ (20)
where the difference PQP-Q is a signed measure on $\mathcal{X}$.
@@
Fix the prior $\pi\in(0,1)$. For generator $g$, let $P_{\rm g}$ denote the distribution of $g(Z)$, when $Z\sim P_{Z}$. Define $D^{(\pi)}$ as
$D^{(\pi)}(g)$ $\triangleq\max_{h}\mathcal{L}_{\text{BG}}^{(\pi)}(g,h),$ (21)
with the maximum taken over all discriminators, i.e., all $h:\mathcal{X}\mapsto[0,1]$. Then, $D^{(\pi)}$ satisfies the following property
$D^{(\pi)}(g)+D^{(1-\pi)}(g)$ $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (22)
with equality when $\pi=0.5$.
@@
From the definition (18), we have
@@
$-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$ (23)
which noting that $g(Z)$ is distributed as $P_{\rm g}$ can be written as
@@
Setting the signed metric $P_{\text{data}}-P_{\rm g}$ to be the difference of densities, we can conclude that
$\Sigma(g)$ $\triangleq\max_{h:\mathcal{X}\mapsto[0,1]}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (25)
$=\max_{h:\mathcal{X}\mapsto[0,1]}\left[\int h(X)\mathrm{d}(P_{\text{data}}-P_{\rm g})\right]$ (26)
@@
$\Sigma(g)\leq\max_{h}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)\right]+\max_{h}\left[\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right],$ (28)
and using the definition of $D^{(\pi)}$ to conclude that
@@
$\leq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (30)
which proves the inequality. We further note that with $\pi=0.5$
@@
$=\max_{h}\left[\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)+\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)\right]$
$=2D^{(0.5)}(g)$ (31)
which indicates that equality is achieved with $\pi=0.5$. ∎
The above result indicates that using BOLT-GAN, the generator distribution converges to the data distribution at least as strong as minimizing the TV between the two distributions.
@@
Let $D^{(\pi)}$ be as in Theorem 3. Assume for a given generator $g$, we have
$\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}\leq\varepsilon$ (32)
@@
$2\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ $\geq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (34)
@@
$\varepsilon$ $\geq\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ (36)
$\;\geq\;\frac{1}{2}\operatorname{TV}\!\bigl(P_{\mathrm{data}},P_{\rm g}\bigr).$ (37)
which concludes the proof. ∎
The above result suggests that the BOLT-GAN objective encourages the generated distribution $P_{g}$ to approach the true data distribution $P_{\mathrm{data}}$ under the total variation (TV) metric.
3.3 Regularized BOLT-GAN via Lipschitz
Lipschitz critics are known to stabilize adversarial training and provide informative, non-saturating gradients Arjovsky et al., (2017); Gulrajani et al., (2017); Miyato et al., (2018). Motivated by this, we enforce $1$-Lipschitz constraint on the discriminator. Intuitively, this can lead to better gradients while training the generator. This can reduce gradient explosion or vanishing, as well as the variance of stochastic updates flowing through $\nabla_{\phi}\,h_{\theta}(g_{\phi}(z))$. In this section, we show that under this modification, the BOLT-GAN implies a weaker notion of convergence and better stability.
Lipschitz BOLT-GAN
The Lipschitz BOLT-GAN trains the generator and discriminator by solving the min-max game (17) , subject to Lipschitz continuity of the discriminator. More precisely, it solves
$\min_{\phi}\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (38)
where $\mathcal{H}_{\mathrm{Lip}}$ is the set of functions $h_{\theta}:\mathcal{X}\mapsto[-1,0]$ and
$\forall\,x,y\in\mathcal{X}\},$
The Lipschitz continuity further imposes a smoothness regularization on the BOLT-GAN discriminator.
For simplicity, we drop the subscript $\theta$ in $h_{\theta}$, stating the following theorem. 4.
Theorem 4 (Convergence of Lipchitz BOLT-GAN).
Fix $0<\pi\leq 0.5$, and define the bounded 1-Lipschitz class of discriminator models as
$\forall\,x,y\in\mathcal{X}\},$
for some norm $\|\cdot\|$ defined on data space $\mathcal{X}$. For the generator model $g$, define
$D^{(\pi)}_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}L^{(\pi)}_{\mathrm{BG}}(g,h),$ (41)
with $L^{(\pi)}_{\mathrm{BG}}(g,h)$ being defined in (18). Then, for any prior $0<\pi\leq 0.5$, we have
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}\big(P_{\mathrm{data}},P_{\rm g}\big).$ (42)
where $W1$ is the Lipchitz Wasserstein‑1 measure defined in (46) below.
Proof.
Starting from (3.2) in the proof of Theorem 3, we can write
$\Sigma_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (43)
$=\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$ (44)
In the supplementary file, we show that for $0<\pi\leq 0.5$,
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (45)
and therefore
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$
The Kantorovich-Rubinstein duality (Villani et al.,, 2009) gives
$W_{1}(P_{\mathrm{data}},P_{g})=$ (46)
$\sup_{h\in\mathcal{H}_{\mathrm{Lip}}^{+}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\},$
where $\mathcal{H}_{\mathrm{Lip}}^{+}$ denotes the set of all 1-Lipschitz functions on $\mathcal{X}$
$\mathcal{H}_{\mathrm{Lip}}^{+}=\{h:|h(x)-h(y)|\leq\|x-y\|\hskip 2.84526pt\forall\,x,y\in\mathcal{X}\}.$ (47)
Noting that $\mathcal{H}_{\mathrm{Lip}}\subset\mathcal{H}_{\mathrm{Lip}}^{+}$, we can write
@@ 3.3 Regularized BOLT-GAN via Lipschitz
$\geq D^{(\pi)}_{\mathrm{Lip}}(g).$ (49)
This concludes the proof. ∎
The above result indicates an interesting dichotomy. While the WGAN implicitly minimizes the Earth-Mover (Wasserstein-1) distance between $P_{\mathrm{data}}$ and $P_{g}$, the BOLT-GAN objective can be viewed as implicitly minimizing the total variation (TV) distance. The Lipschitz BOLT-GAN, on the other hand, bridges these two perspectives by enforcing Lipschitz continuity, thereby promoting a smoother optimization landscape. This intuition may be suggestive of the relative advantages of these three methods.
Lipschitz enforcement. We impose Lipschitzness on $h_{\theta}$ by adding a gradient penalty (GP, $\lambda_{\mathrm{GP}}{=}10$) (Gulrajani et al.,, 2017): for $\hat{x}=\epsilon x+(1-\epsilon)g_{\phi}(z)$ with $\epsilon\!\sim\!\mathcal{U}(0,1)$, we add $\lambda_{\mathrm{GP}}\,\mathbb{E}_{\hat{x}}\!\big(\|\nabla_{\hat{x}}h_{\theta}(\hat{x})\|_{2}-1\big)^{2}$ to the loss. Alternatives (spectral normalization, weight clipping) and implementation details are discussed in the supplement.

### v2

@@ 3.1 Bayes Optimal Discriminator via BOLT
See Appendix B.∎
Theorem 2 further enables us to express the MAP classifier in terms of the likelihood ratio and rewrite the Bayes error as a hinge expectation.
Corollary 1 (Optimal BOLT function and hinge form of $\varepsilon_{\rm Bayes}$).
Assume $q_{1},q_{2}>0$ and that $P_{1},P_{2}$ admit densities $p_{1},p_{2}$ w.r.t. a common dominating measure. Define the likelihood ratio $U(x):=p_{1}(x)/p_{2}(x)$ on $\{p_{2}(x)>0\}$ and the prior threshold $\tau:=q_{2}/q_{1}$. Then a pointwise optimizer of the BOLT functional over $h:\mathcal{X}\to[0,1]$ is
$h^{\star}(x)\;=\;\begin{cases}0,&U(x)<\tau,\\[2.0pt] 1,&U(x)\geq\tau,\end{cases}$ (15)
and, defining the hinge map $t_{0}(u):=[\,q_{2}-q_{1}u\,]_{+}$, one has
$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}}\;=\;q_{2}-\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))\big].$ (16)
Proof.
For $x$ with $p_{2}(x)>0$, Bayes’ rule gives
$\eta(x):=\Pr(C=1\mid X=x)=\frac{q_{1}U(x)}{q_{1}U(x)+q_{2}}.$ (17)
Hence, $\eta(x)\gtrless\tfrac{1}{2}$ iff $U(x)\gtrless\tau$ with $\tau:=q_{2}/q_{1}$ . Substituting this relation into Theorem 2 yields (15). The hinge representation (16) follows by rewriting the tight BOLT bound at $h^{\star}$ in terms of $U$. ∎
Corollary 1 shows that $\varepsilon_{\rm Bayes}$ can be written as an expectation of a hinge of $U$. In practice, a discriminator trained via empirical BOLT risk can be used to form a finite-sample likelihood-ratio estimate $\widehat{U}$ (see Appendix B). Plugging $\widehat{U}$ into Corollary 1, we can estimate the MAP classifier and BER. The following result characterizes the bias and variance of the BER estimator.
Theorem 3 (BER estimator).
Let $\widehat{U}:\mathcal{X}\to\mathbb{R}$ be an estimator of the likelihood ratio $U$, and let $\widehat{q}_{1},\widehat{q}_{2}\in(0,1)$ be empirical class priors computed on M samples. Define $\widehat{\varepsilon}_{\rm BOLT}$ as the plug-in estimator obtained from the hinge representation in Corollary 1 by sample averaging with $M$ samples. Assume $\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\leq\varepsilon_{0}$. Then, we have
$\big|\mathbb{E}[\widehat{\varepsilon}_{\rm BOLT}]-\varepsilon_{\rm Bayes}\big|\leq q_{1}\,\varepsilon_{0}+O(M^{-1/2}),$ (18)
and
$\mathrm{Var}(\widehat{\varepsilon}_{\rm BOLT})=O(M^{-1}).$ (19)
Proof.
See Appendix B.∎
3.2 Generation with Maximum BER
The results of Section 3.1 allow us to interpret a BOLT-trained discriminator as an estimator of the Bayes-optimal classifier for the real-fake classification task, and to view its objective as a tight surrogate for the discrimination BER. We therefore construct an adversarial objective where, for a fixed generator $g$, the discriminator $h$ is trained via the (population) BOLT risk under prior $\pi$, and the generator is trained to maximize the induced BER (equivalently, minimize the discriminator best achievable performance). Intuitively, this pushes $P_{\rm g}$ toward $P_{\rm data}$ by making real and fake samples maximally confusable. We characterize this behavior by showing that the resulting maximum BER problem minimizes the TV between $P_{\rm data}$ and $P_{\rm g}$.
Maximum BER objective
Consider a GAN with generator $g$ and a bounded discriminator $h:\mathcal{X}\to[0,1]$. Let $P_{\rm g}$ denote the distribution of $g(Z)$ for $Z\sim P_{Z}$, and define $\pi=\Pr\{C=1\}$, i.e. probability of sampling a real data-point. We train $h$ by the empirical BOLT risk, and train $g$ adversarially to maximize the induced discrimination BER. Following the BOLT formulation, the discrimination BER is estimated (up to a constant) via
$\max_{h}\;\mathcal{L}_{\rm BG}^{(\pi)}(g,h),$ (20)
where $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ is defined as [^3] Note that $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ is the negative of the (population) BOLT risk, up to the additive constant $\pi$.
$\mathcal{L}_{\rm BG}^{(\pi)}(g,h)\;:=\;$ $\pi\,\mathbb{E}_{X\sim P_{\rm data}}\!\big[h(X)\big]$
$-\,(1-\pi)\,\mathbb{E}_{Z\sim P_{Z}}\!\big[h(g(Z))\big].$ (21)
By Theorem 2, thresholding the BOLT-optimal discriminator yields the Bayes optimal discrimination rule. Accordingly, the generator can adversarially fool the Bayes optimal discrimination by solving the Max-BER problem that is given by
$\min_{g}\max_{h}\;\mathcal{L}_{\text{BG}}^{(\pi)}(g,h).$ (22)
Generator distribution
@@ 3.2 Generation with Maximum BER
Intuitively, the Max-BER problem leads to learning data distribution. We next show this and characterize the metric, under which the generator distribution $P_{\rm g}$ converges to $P_{\rm data}$. To this end, let us first define total variation (TV).
@@
The total variation between two distributions $P$ and $Q$ defined on a measurable space (X, B) is defined as
$\mathrm{TV}(P,Q)=\sup_{A\in\cal{B}}|P(A)-Q(A)|,$ (23)
where $\mathcal{B}$ is the $\sigma$-algebra on $\mathcal{X}$.
In the sequel, we use the standard variational representation of TV given in the following lemma.
Lemma 1 (Dudley (2002a)).
@@
$\mathrm{TV}(P,Q)=\sup_{f:\mathcal{X}\to[0,1]}\int f\,\mathrm{d}(P-Q),$ (24)
where the difference PQP-Q is a signed measure on $\mathcal{B}$.
@@
Fix the prior $\pi\in(0,1)$. For generator $g$, let $P_{\rm g}$ denote the distribution of $g(Z)$, when $Z\sim P_{Z}$. Define $\mathcal{D}^{(\pi)}$ as
$\mathcal{D}^{(\pi)}(g)$ $\triangleq\sup_{h}\mathcal{L}_{\text{BG}}^{(\pi)}(g,h),$ (25)
with the maximum taken over all bounded discriminators, i.e., all $h:\mathcal{X}\mapsto[0,1]$. Then, $\mathcal{D}^{(\pi)}$ satisfies
$\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g)$ $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (26)
with equality holding when $\pi=0.5$.
@@
From the definition (3.2), we have
@@
$-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))].$ (27)
Noting that $g(Z)$ has distribution $P_{\rm g}$ , we can write
@@
Then
$\Sigma(g)$ $\triangleq\sup_{h:\mathcal{X}\mapsto[0,1]}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (29)
$=\sup_{h:\mathcal{X}\mapsto[0,1]}\left[\int h(X)\mathrm{d}(P_{\text{data}}-P_{\rm g})\right]$ (30)
@@
$\Sigma(g)\leq\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)\right]+\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right],$ (32)
and the definition of $\mathcal{D}^{(\pi)}$ to conclude that
@@
$\leq\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g).$ (34)
This proves the inequality. We next note that at $\pi=0.5$
@@
$=\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)+\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)\right]$
$=2\mathcal{D}^{(0.5)}(g).$ (35)
This concludes the proof. ∎
Theorem 4 indicates that the solution to the Max-BER problem leads to $P_{\rm g}$ converging $P_{\rm data}$ in TV. In particular, under the balanced prior $\pi=0.5$, we obtain the exact identity $\mathrm{TV}(P_{\rm data},P_{\rm g})=2\mathcal{D}^{(0.5)}(g)$, so the Max-BER problem is equivalent to minimizing TV. For general $\pi\in(0,1)$, Theorem 4 relates TV to the pair of games specified by $\mathcal{D}^{(\pi)}(g)$ and $\mathcal{D}^{(1-\pi)}(g)$, and hence the convergence in TV is guaranteed when the worst-case objective converges to zero.
@@
Let $\mathcal{D}^{(\pi)}$ be defined as in Theorem 4. Assume for a given generator $g$, we have
$\max\left\{\mathcal{D}^{(\pi)}(g),\mathcal{D}^{(1-\pi)}(g)\right\}\leq\varepsilon.$ (36)
@@
$2\max\left\{\mathcal{D}^{(\pi)}(g),\mathcal{D}^{(1-\pi)}(g)\right\}$ $\geq\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g)$
@@
This concludes the proof. ∎
4 BOLT-GAN Framework
With an unregularized bounded-discriminator, i.e., the Max-BER setting, the optimal discriminator may be highly non-smooth, and the training of a parameterized model can be unstable. Theorem 4 shows that under balanced priors this problem is directly linked to TV, a strong discrepancy that can yield poorly behaved gradients in adversarial optimization. For stable training, we propose BOLT-GAN, which solves the Max-BER problem while restricting the discriminator to be 1-Lipschitz.
More precisely, BOLT-GAN solves the following constrained Max-BER problem
$\min_{g}\max_{h\in\mathcal{H}_{\rm Lip}}\;\mathcal{L}^{(\pi)}_{\rm BG}(g,h),$ (39)
where $\mathcal{H}_{\mathrm{Lip}}$ denotes the class of bounded 1-Lipschitz discriminators, i.e.,
$\forall\,x,y\in\mathcal{X}\}.$
We next characterize the convergence of BOLT-GAN.
4.1 BOLT-GAN Generator Distribution
Theorem 5 shows that the constrained Max-BER problem solved by BOLT-GAN minimizes an induced IPM that is upper-bounded by the Wasserstein-1 distance between $P_{\rm g}$ and $P_{\rm data}$. Before stating the main result, let us define the Wasserstein-1 distance.
Definition 2 (Wasserstein-1 distance).
Let $(\mathcal{X},\|\cdot\|)$ be a metric space and let $P,Q$ be probability measures on $\mathcal{X}$ with finite first moment. The Wasserstein-1 distance is
$W_{1}(P,Q)\triangleq\inf_{\gamma\in\Pi(P,Q)}\mathbb{E}_{(X,Y)\sim\gamma}\big[\|X-Y\|\big],$ (41)
where $\Pi(P,Q)$ is the set of couplings of $P$ and $Q$.
Computing the Wasserstein-1 distance between two general probability measures is intractable in closed form. The Kantorovich-Rubinstein duality result Villani (2009) gives an alternative formulation which enables estimating this distance by sampling.
Lemma 2 (Kantorovich-Rubinstein duality Villani (2009)).
Let $P,Q$ be probability measures on $(\mathcal{X},\|\cdot\|)$ with finite first moment. Then
$W_{1}(P,Q)=\sup_{f:\,\mathrm{Lip}(f)\leq 1}\left\{\mathbb{E}_{X\sim P}[f(X)]-\mathbb{E}_{X\sim Q}[f(X)]\right\}.$
We next present the Theorem 5, which characterizes the converging behavior of the generator when trained by BOLT-GAN.
Theorem 5 (BOLT-GAN Convergence).
Fix $0<\pi\leq 0.5$, for the generator $g$, define
$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathcal{L}_{\rm BG}^{(\pi)}(g,h),$ (42)
with $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ defined in (3.2). Then, for any prior $0<\pi\leq 0.5$, we have
$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}\big(P_{\mathrm{data}},P_{\rm g}\big).$ (43)
where $W_{1}$ denotes the Wasserstein–$1$ distance.
Proof.
Starting with (3.2) in the proof of Theorem 4, we can write
$\Sigma_{\mathrm{Lip}}(g)\triangleq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (44)
$=\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$ (45)
In the Appendix C, we show that for $0<\pi\leq 0.5$,
$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq\Sigma_{\mathrm{Lip}}(g).$ (46)
@@ 4.1 BOLT-GAN Generator Distribution
$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$
Let $\mathcal{H}_{\mathrm{Lip}}^{+}$ denotes the set of 1-Lipschitz functions, i.e.,
$\mathcal{H}_{\mathrm{Lip}}^{+}=\{h:|h(x)-h(y)|\leq\|x-y\|\hskip 2.84526pt\forall\,x,y\in\mathcal{X}\}.$
Noting that $\mathcal{H}_{\mathrm{Lip}}\subset\mathcal{H}_{\mathrm{Lip}}^{+}$, we can use the Kantorovich-Rubinstein duality to write
@@
$\geq\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g),$ (48)
which concludes the proof. ∎
We note that $\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)$ is an IPM between $P_{g}$ and $P_{\rm{data}}$, and it is minimized by BOLT-GAN as indicated in (39). Theorem 5 implies that convergence in the Wasserstein-1 distance suffices to guarantee convergence in the BOLT-GAN IPM, while the converse does not necessarily hold. Our numerical experiments further show that this weaker distance metric leads to more stable training.
4.2 Computational Lipschitz Enforcement
In practice, we parameterize the discriminator as $h_{\theta}(X)=\sigma(f_{\theta}(x))$ for some score function $f_{\theta}(x)$, where $\sigma$ denotes the sigmoid function. The 1-Lipschitz constraint is computationally enforced on the score $f_{\theta}$ using a gradient penalty (GP) (Gulrajani et al., 2017): for $\hat{X}=\varepsilon X+(1-\varepsilon)g(Z)$ with $\varepsilon\sim{\rm Unif}(0,1)$, we add
$\lambda_{\rm GP}\,\mathbb{E}\Bigl(\|\nabla_{\hat{X}}f_{\theta}(\hat{X})\|_{2}-1\Bigr)^{2}$ (49)
to the discriminator loss for some regularizer $\lambda_{\rm GP}$. Since the sigmoid function $\sigma$ is Lipschitz, this enforcement implies that $h$ is also Lipschitz. [^4] Indeed, $h$ is Lipschitz with $\mathrm{Lip}(h)\leq\mathrm{Lip}(\sigma)\,\mathrm{Lip}(f_{\theta})$.